# Primordial Features in light of the Effective Field Theory of Large-Scale Structure

### v1

@@ 4 Global features: linear and logarithmic oscillations
Beyond theoretical motivations, the study of linear and logarithmic oscillations superimposed on a power-law spectrum offers valuable pedagogical insights. These oscillatory features have been extensively explored in the literature due to their simplicity and theoretical relevance (see e.g., Refs. [7, 8] in the context of CMB analyses, and Refs. [38, 39, 125, 41] in the context of LSS analyses). Here, following Refs. [126] and to remain model agnostic, we perform the usual (template-based) search for features [7, 8]
@@ 5.2 Damped logarithmic oscillations
$\mathcal{P}_{\zeta}(k)=\mathcal{P}_{\zeta,0}(k)\left[1+A_{\rm log}\sin\left(\omega_{\rm log}~{}\frac{k}{k_{*}}+\phi_{\rm log}\right)e^{-\frac{\beta^{2}(k-\mu)^{2}}{2k_{*}^{2}}}\right]~{},$ (5.2)
where $\mu$ and $\beta$ are respectively the center and the width of the Gaussian envelope. In this model, the phase is determined by $\mu$ and $\beta$ in such a way that the peak of the Gaussian is located at a maximum of the oscillations: $\phi_{\rm log}=-\omega_{\rm log}\ln(\mu/k_{*})$. Let us note that when $\beta\to 0$, we recover the global logarithmic oscillations explored in Section 4. Following Ref. [148], we impose the following priors: $A_{\rm log}\in[0,0.5]$, $\log_{10}(\omega_{\rm log})\in[1.0,1.8]$, $\mu\in[0.001,0.175]\ {\rm Mpc}^{-1}$ and $\beta\in[0,10]$. According to the methodology developed in Section 4.1, we can safely use these priors in the LSS analysis. In Fig. 12, we plot the non-linear matter power spectrum residuals for the effective redshifts considered in our analysis. We further note that for large $\beta$, i.e. very localized oscillatory features, the constraints on the amplitude $A_{\rm log}$ can be significantly relaxed with respect to the global logarithmic oscillations, as for the choice of cosmological parameters shown in Fig. 4.
@@
$A_{\rm log}^{\rm LSS}$ $<0.314\,,$
$A_{\rm log}^{\rm CMB}$ $<0.237\,.$
Once again, we can see that LSS constraints are competitive and complementary to the CMB constraints, and offer a promising way of tightly constraining the amplitude of primordial features with the next generation of LSS surveys.
@@ 7 Conclusions
In this work, we applied the effective field theory of large-scale structure (EFTofLSS) to analyze the galaxy power spectrum multipoles measured by BOSS LRG and eBOSS QSO, in combination with CMB (T&E) data, to constrain some representative primordial features that cover a wide range of inflationary dynamics. A key strength of our approach is its ability to incorporate not only the non-linear effects from the long-wavelength modes (through IR-resummation) but also to account for small-scale non-linearities thanks to the one-loop EFT contributions, allowing us to analyze scales up to $k=0.24h{\rm Mpc}^{-1}$. In addition, we emphasize that our IR-resummation scheme does not rely on the wiggle/no-wiggle split procedure (as done in past literature [38, 41, 125]), but on an analytical formula (implemented in PyBird) which is (in principle) independent of the primordial feature scenario considered. We summarize our results as follows:
@@
There are several ways in which our work can be improved and extended. First, the frequency range of the global primordial features can be expanded to include higher frequencies by reducing the bandwidth $\Delta k$, following the approach in Refs. [38, 41]. Second, our analysis can be extended beyond two-point statistics by incorporating bispectrum measurements, as recently developed in Refs. [179, 180], which provide additional sensitivity to primordial features. Third, a joint reconstruction of the PPS using both CMB and LSS data could be performed, for instance, by applying the regularized modified Richardson-Lucy methodology simultaneously to both datasets. Finally, our analysis pipeline is readily applicable to power spectrum measurements from DESI [46] and the forthcoming Euclid mission [48, 49], which will provide higher-precision data and further refine constraints on primordial features.
@@ Acknowledgments
We thank Xingang Chen, Guido D’Amico, Fabio Finelli, Vivian Poulin and Pierre Zhang for insightful feedback and comments at various stages of the project, as well as Wuhyun Sohn for interesting discussions and for sharing the regMRL reconstructed PPS. We would also like to acknowledge all the participants to the ModIC 2024 workshop for the fruitful interdisciplinary exchange of ideas. R.C. is funded by the Czech Ministry of Education, Youth and Sports (MEYS) and European Structural and Investment Funds (ESIF) under project number CZ.02.01.01/00/22_008/0004632. A.S. would like to acknowledge the support by National Research Foundation of Korea 2021M3F7A1082056, and the support of the Korea Institute for Advanced Study (KIAS) grant funded by the government of Korea. DKH acknowledges India-Italy “RELIC - Reconstructing Early and Late events In Cosmology” mobility program and the Indo-French Centre for the Promotion of Advanced Research – CEFIPRA grant no. 6704-4. This work was supported by the computational resources from the LUPM’s cloud computing infrastructure founded by Ocevu labex and France-Grilles, as well as the high-performance computing cluster Seondeok at the Korea Astronomy and Space Science Institute. We acknowledge the use of the Boltzmann solver class [181, 182] for the computation of theoretical observables, Montepython [53, 54] for the sampling and GetDist [75] for the post-processing of our results. We also acknowledge the use of the standard python libraries for scientific computing, such as numpy [183], scipy [184] and matplotlib [185].

### v2

@@ 4 Global features: linear and logarithmic oscillations
Beyond theoretical motivations, the study of linear and logarithmic oscillations superimposed on a power-law spectrum offers valuable pedagogical insights. These oscillatory features have been extensively explored in the literature due to their simplicity and theoretical relevance (see e.g., Refs. [7, 8] in the context of CMB analyses, and Refs. [38, 39, 125, 41] in the context of LSS analyses). Here, following Ref. [126] and to remain model agnostic, we perform the usual (template-based) search for features [7, 8]
@@ 5.2 Damped logarithmic oscillations
$\mathcal{P}_{\zeta}(k)=\mathcal{P}_{\zeta,0}(k)\left[1+A_{\rm log}\cos\left(\omega_{\rm log}~{}\frac{k}{k_{*}}+\phi_{\rm log}\right)e^{-\frac{\beta^{2}(k-\mu)^{2}}{2k_{*}^{2}}}\right]~{},$ (5.2)
where $\mu$ and $\beta$ are respectively the center and the width of the Gaussian envelope. In this model, the phase is determined by $\mu$ and $\beta$ in such a way that the peak of the Gaussian is located at a maximum of the oscillations: $\phi_{\rm log}=-\omega_{\rm log}\ln(\mu/k_{*})$. Let us note that when $\beta\to 0$, we recover the global logarithmic oscillations explored in Section 4. Following Ref. [148], we impose the following priors: $A_{\rm log}\in[0,0.5]$, $\log_{10}(\omega_{\rm log})\in[1.0,1.8]$, $\mu\in[0.001,0.175]\ {\rm Mpc}^{-1}$ and $\beta\in[0,10]$. According to the methodology developed in Section 4.1, we can safely use these priors in the LSS analysis. In Fig. 12, we plot the non-linear matter power spectrum residuals for the effective redshifts considered in our analysis. We note that for large $\beta$, i.e., very localized oscillatory features, the constraints on the amplitude $A_{\rm log}$ can be significantly relaxed with respect to the global logarithmic oscillations, as for the choice of cosmological parameters shown in Fig. 12.
@@
$A_{\rm log}^{\rm LSS}$ $<0.262\,,$
$A_{\rm log}^{\rm CMB}$ $<0.237\,,$
$A_{\rm log}^{\rm CMB+LSS}$ $<0.124\,.$
Once again, we can see that LSS constraints alone are competitive and complementary to the CMB constraints, and offer a promising way of tightly constraining the amplitude of primordial features with the next generation of LSS surveys.
@@ 7 Conclusions
In this work, we applied the effective field theory of large-scale structure (EFTofLSS) to analyze the galaxy power spectrum multipoles measured by BOSS LRG and eBOSS QSO, in combination with CMB (T&E) data, to constrain some representative primordial features that cover a wide range of inflationary dynamics. A key strength of our approach is its ability to incorporate not only the non-linear effects from the long-wavelength modes (through IR-resummation) but also to account for small-scale non-linearities thanks to the one-loop EFT contributions, allowing us to analyze scales up to $k=0.24h{\rm Mpc}^{-1}$. In addition, we emphasize that our IR-resummation scheme does not rely on the wiggle/no-wiggle split procedure (as done in past literature [38, 41, 125]), but on an analytical formula (implemented in PyBird) which is (in principle) independent of the primordial feature scenario considered (see also [179] for a similar LPT-based approach). We summarize our results as follows:
@@
There are several ways in which our work can be improved and extended. First, the frequency range of the global primordial features can be expanded to include higher frequencies by reducing the bandwidth $\Delta k$, following the approach in Refs. [38, 41]. Second, our analysis can be extended beyond two-point statistics by incorporating bispectrum measurements, as recently developed in Refs. [180, 181, 179], which provide additional sensitivity to primordial features. Third, a joint reconstruction of the PPS using both CMB and LSS data could be performed, for instance, by applying the regularized modified Richardson-Lucy methodology simultaneously to both datasets. Finally, our analysis pipeline is readily applicable to power spectrum measurements from DESI [46] and the forthcoming Euclid mission [48, 49], which will provide higher-precision data and further refine constraints on primordial features.
@@ Acknowledgments
We thank Xingang Chen, Guido D’Amico, Fabio Finelli, Thiago Mergulhão, Vivian Poulin, and Pierre Zhang for insightful feedback and comments at various stages of the project, as well as Wuhyun Sohn for interesting discussions and for sharing the regMRL reconstructed PPS. We would also like to acknowledge all the participants to the ModIC 2024 workshop for the fruitful interdisciplinary exchange of ideas. R.C. is funded by the Czech Ministry of Education, Youth and Sports (MEYS) and European Structural and Investment Funds (ESIF) under project number CZ.02.01.01/00/22_008/0004632. A.S. would like to acknowledge the support by National Research Foundation of Korea 2021M3F7A1082056, and the support of the Korea Institute for Advanced Study (KIAS) grant funded by the government of Korea. DKH acknowledges India-Italy “RELIC - Reconstructing Early and Late events In Cosmology” mobility program and the Indo-French Centre for the Promotion of Advanced Research – CEFIPRA grant no. 6704-4. TS acknowledge the European Union’s Horizon Europe research and innovation programme under the Marie Skłodowska-Curie Staff Exchange grant agreement No 101086085 – ASYMMETRY. This work was supported by the computational resources from the LUPM’s cloud computing infrastructure founded by Ocevu labex and France-Grilles, as well as the high-performance computing cluster Seondeok at the Korea Astronomy and Space Science Institute. We acknowledge the use of the Boltzmann solver class [182, 183] for the computation of theoretical observables, Montepython [53, 54] for the sampling and GetDist [75] for the post-processing of our results. We also acknowledge the use of the standard python libraries for scientific computing, such as numpy [184], scipy [185] and matplotlib [186].